There are two types of borrowers in the credit market for motorcycle loans, type X, and type R. They both need a loan for $37,600, for a Ducati. Type X repays 96% of the time and type R repays 84% of the time. If the lending bank, Street Bank, has full information, then the risk premium for X is ____% and for R it is ____%.

Group of answer choices

A) 5.1; 19

B) 4.2; 19

C) 5.1; 15.5

D) 4.2; 15.5
Answer #1

Risk premium = Non payment risk%/ Repayment%

X repays 96% , which means non payment risk % = 4%

Y repays 84%, which means non payment risk % = 16%

Risk premium of X = 4/96 = 4.2

Risk premium of Y = 16/84 = 19

Therefore the correct option is B